# Relativistic and Electron Correlation Effects in Static Dipole Polarizabilities for Main-Group Elements

### v1

@@ 4.8.1 Comparison with literature
Ref. 3 --- - --- - $1.38$ $2.66$ $11.08$ $16.78\pm 0.02$ $27.32\pm 0.2$ $35.0\pm 2$ $58.0\pm 6$
@@
The DC results for the heavy atoms, including krypton (Kr, $Z=36$), xenon (Xe, $Z=54$), radon (Rn, $Z=86$), and Og, show excellent agreement with the reference values in Ref. 3. For the lighter atoms, like Ne, and argon (Ar, $Z=18$), the polarizabilities calculated in this study are slightly lower than the recommended values as shown in Table 8. Moreover, the NR, SR, and DC values are 2.57 for Ne in this work, which is lower than experimental 113, 114, 2, 115 and theoretical values,116, 57, 117, 85, 118, 119 i.e., 2.66. Additionally, the SR and DC values are 11.02 for Ar in this work, which is lower than experimental 120, 121, 114, 57, 115 and theoretical values,57, 88, 92, 122 i.e., 11.08.
@@
In addition, the most accurate theoretical and experimental values for Group 18 elements are collected in Table S32 to validate the results in this work. Next, a detailed comparison between this work and results in the literature is presented. Only computational and experimental values close to the recommended values in Ref. 3 are compared. Only values computed with CCSD(T) methods in the literature are considered.
The NR, SR, and DC values are 1.38 for He in this work, which align well with the experimental values 120, 123, 1, 115 and other theoretical values.124, 125, 115, 126 The deviation is mainly attributed to the $\Delta_{\text{others}}$, including Breit and QED effects, among others.
The relativistic effects on dipole polarizabilities for Kr are small. Therefore, all NR, SR, and DC values (16.78, 16.77, and 16.80, respectively) for Kr are in good agreement with the experimental values ($16.766\pm 0.008$,121 $16.782\pm 0.005$,113 $16.740$,114, 127 and $16.734$ 114). The SR value also agrees well with the theoretical values ($16.80\pm 0.13$ 127 and 16.736 128) obtained using scalar-relativistic CCSD(T) methods in the literature. The DC value is also close to the recent MBPT value (16.800) including the Gaunt contribution.82
The DC value (27.20) for Xe is in good agreement with the experimental values (27.34227.34227.34227.342129 $27.078\pm 0.050$,113 and $27.292$ 114). The SR value (27.01) is also close to other computational values ($27.06\pm 0.27$,130 $27.2937\pm 0.0003$,116 $27.42$,131 $28.4\pm 0.5$,132 and $27.508$ 133).
There is no experimental data available for comparison for Rn in the literature. The DC value (35.07) for Rn is in excellent agreement with other computational DC results ($35.04\pm 1.8$,88, 9 $35.0$,134 $35.3$,135 and $35.53\pm 0.36$ 136) in the literature.
There is no experimental data available for comparison for Og in the literature. The DC value (58.24) for Og in this work is in excellent agreement with other computational DC results ($57.98$ 137 and $56.5\pm 1.8$ 136). Moreover, it is observed that the hyperpolarizability using Eq. (14) for Og is negative in this work. Therefore, the values fitted using Eq. (15) are also computed, and the NR, SR, and DC values are 45.28, 43.77, and $57.90\pm 0.06$, respectively, as shown in Table S24.
Jerabek et al. reported NR, SR, and DC values for Og, using the same method and basis set as in this work, as 45.30, 43.78, and 57.98, respectively. The Gaunt contribution is also included in the DC value. In Ref. 137, the numerical differentiation method is used, and only $\alpha$ is fitted.138 It can be seen that the NR and SR values fitted by Eq. (15) perfectly align with the counterparts reported in Ref. 137. The difference in DC values (0.08) could be mainly attributed to the Gaunt contribution.
Additionally, the 4C Hamiltonian is used in this work, while a 2C Hamiltonian is used in Ref. 137. To study the difference due to the 2C and 4C Hamiltonians used, three test calculations were performed. The first and second calculations used a 2C Hamiltonian with 18 and 40 electrons correlated, denoted as 2C-DC-CC@dyall.acv4z@(core 18)[vir 278] and 2C-DC-CC@dyall.acv4z@(core 40)[vir 278], respectively, as shown in Table S16. The CCSD(T) results fitted using Eq. (15) are 57.86 and 57.96, respectively. The DHF values for both cases are 60.58. The third test calculation used 2C-CC with 50 electrons correlated, and the Gaunt contribution to $\alpha$ was also evaluated at the DHF level, denoted as 2C-DC-Gaunt-CC@dyall.acv4z@(core 50)[vir 278] in Table S16. The DHF and CCSD(T) results fitted using Eq. (15) are 60.45 and 57.79, respectively. The difference between the CCSD(T) results of the third test and the CCSD(T) value reported in Ref. 137 is about 0.19 a.u., which could be attributed to the Gaunt contribution evaluated at the CCSD(T) level in Ref. 137. Comparing the results between the second and third calculations, the Gaunt contribution to $\alpha$ at the DHF level is about $-0.13$. Therefore, the total Gaunt contribution is about 0.06, which is close to the observed difference (0.08). These tests suggest that the difference between the 2C and 4C methods is negligible, and the Gaunt contribution to $\alpha$ is also small.
4.8.2 Uncertainty estimation
Next, the uncertainty excluding $\Delta P_{\text{fitting}}$ is estimated for all Group 18 elements. For He, Ne, and Ar, accurate results are available in the literature. Therefore, the deviation between the DC value and the recommended value in Ref. 3 is taken as the uncertainty, which is mainly due to $\Delta P_{\text{others}}$.
Several SR test calculations with the d-aug-dyall.acv4z basis set are performed for Kr, Xe, Rn, and Og. Compared to the SR CCSD(T) values obtained with dyall.acv4z, the differences are about 0.01, 0.01, 0.01, and 0.13, respectively. The difference for Og with $\alpha$ obtained using Eq. (15) is 0.12. These values are used as the error due to the finite basis set ($\Delta P_{\text{basis}}$).
The difference between 2C and 4C Hamiltonians is studied for Ne, Ar, Xe, and Rn, as shown in Table S8, where the X2C methods with the same number of correlated electrons as in 4C-CC methods are performed. The differences are 0.00, 0.01, 0.01, and 0.03 a.u., respectively. The difference for Ar and Xe is 0.01; thus, it is assumed that it is also small for Kr in this work.
The Gaunt contribution has been investigated in Ref. 82, where Gaunt contributions to $\alpha$ are 0.0266, 0.0585, and 0.1813 for Kr, Xe, and Rn, respectively. It can be seen that the error due to 2C and 4C Hamiltonians is negligible. Therefore, only the Gaunt contribution is considered the main error due to $\Delta P_{\text{others}}$ for Kr, Xe, and Rn. For Og, the sum of the difference (0.33) due to using fitting equations defined in Eqs. (14) and (15) and the Gaunt contribution (0.08) is taken as the error due to $\Delta P_{\text{others}}$.
The error due to $\Delta P_{\text{core}}$ and $\Delta P_{\text{vir}}$ is also studied for Kr, Xe, Rn, and Og. For Kr, a DC-CC calculation with 26 correlated electrons was performed compared to the most accurate calculation with 36 correlated electrons in this work; the differences at the CCSD and CCSD(T) levels are both 0.02.
For Xe, a 2C-DC calculation with 8 correlated electrons was performed compared to a 2C-DC calculation with 50 correlated electrons; the differences at the CCSD and CCSD(T) levels are 0.43 and 0.40, respectively. It should be noted that this difference is close to the difference (0.35) between the DHF and CCSD(T) values in the 2C-DC calculation with 50 correlated electrons. This implies that 50 correlated electrons are sufficient for $\alpha$ calculations for Xe.
Similarly, for Rn, a 2C-DC calculation with 18 correlated electrons was performed compared to another 2C-DC calculation with 50 correlated electrons. The difference is about 0.01, suggesting that 50 correlated electrons are sufficient for Rn.
For Og, a difference of 0.43 is obtained when two 2C-DC CC calculations with 18 and 40 correlated electrons were performed. However, this difference is reduced to 0.01 when comparing two DC-CC calculations with 40 and 50 correlated electrons.
Moreover, the difference ($\Delta P_{\text{(T)}}$) between CCSD and CCSD(T) values in the DC-CC calculation is taken as the error due to $\Delta P_{\text{(T)}}$ and truncating higher-order correlation. They are 0.14, 0.20, 0.26, and 0.06 for Kr, Xe, Rn, and Og, respectively. The value for Og with the $\alpha$ obtained by Eq. (15) is 0.04. In conclusion, the errors of 0.14, 0.21, 0.32, and 0.43 for Kr, Xe, Rn, and Og, respectively, are acceptable. The recommended values, obtained from the most accurate calculations with total uncertainty, including the corresponding $\Delta P_{\text{fitting}}$, are listed in Table 8.
4.8.3 The correlation and relativistic effects on polarizabilities
@@
Figure 8: Polarizabilities (in a.u.) of Group 18 elements. (a) Comparison of non-relativistic (NR), scalar-relativistic (SR), and Dirac-Coulomb (DC) dipole polarizabilities. (b) Illustration of the influence of relativistic effects, including SR, spin-orbit coupling (SOC), and DC, on dipole polarizabilities. (c) Examination of the impact of electron correlation on dipole polarizabilities in the presence of various relativistic effects.
@@ 5 Summary
    The atomic dipole polarizabilities obtained by fitting Eq. (15) are also included. The summary tables for the most accurate results for each group are provided.
@@
    The reference atomic dipole polarizabilities from literature, reported in Refs. 3, 13, are sorted by publication year for each group.

### v2

@@ 4 Results
Ref. 1 --- - --- - $1.38$ $2.66$ $11.08$ $16.78\pm 0.02$ $27.32\pm 0.2$ $35.0\pm 2$ $58.0\pm 6$
4.1 Group 1 elements
As shown in LABEL:tab:dipole_summary, the DC values show good agreement with the recommended values from Ref. 1 for rubidium (Rb, $Z=37$) and Fr, but not for lithium (Li, $Z=3$), sodium (Na, $Z=11$), K, and Cs.
For Li, the DC result (164.31) aligns closely with experimental measurements ($164.0\pm 3.4$ 55 and $164.2\pm 1.1$ 56). The recommended value ($164.1125\pm 0.0005$) in Ref. 1 is based on the value reported in Ref. 57. The small deviation of only 0.12% from the recommended value may stem from the exclusion of relativistic effects such as Breit interactions and QED.
For Na, the DC result ($163.27\pm 0.09$) is in close agreement with the experimental value ($161\pm 7.5$) reported by Ma et al.58 Additionally, it slightly exceeds the recommended computational value (162.88) from Refs. 59, 60, which were obtained using a composite scheme developed by Thakkar and Lupinetti. The NR and SR values for Na, calculated as 163.98 and 162.99 respectively, match well with the corresponding computational values of 163.90 and 162.90 reported in Refs. 59, 60. The difference between the DC value in this work and the final recommended value from Refs. 59, 60 can be attributed to the method used to account for the SOC contribution to $\alpha$. In Refs. 59, 60, a value of $-0.02$ a.u. from Ref. 61, calculated at the DHF level, was added to the SR value. In contrast, the SOC contribution in this work was evaluated at the CCSD(T) level as $+0.28$ a.u. It is noteworthy that the SOC contributions at the DHF and CCSD levels in this study are $-0.04$ and $-0.03$ a.u., respectively. Additionally, this study confirms that the NR value is consistent with the recent non-relativistic CCSD(T) value (163.9) reported by Smiałkowski and Tomza at the CCSD(T)/aug-cc-pwCV5Z level,62 further supporting the reliability of the method employed.
For K, the lower bound of the DC value ($290.64\pm 0.24$) is marginally higher than the upper bound of the recommended value ($289.7\pm 0.3$),1 which is based on experimental data.63, 64 However, the DC value obtained using CCSD(T) from Eq. (15) is $289.40\pm 0.20$ (as shown in Table S9), in good agreement with the recommended value from Ref. 1. The slight discrepancy may arise from the invalidity of a negative $\gamma$ parameter obtained from Eq. (14), particularly when $\tau>0.1\%$ is used in Eq. (16).
For Cs, the DC value ($398.24\pm 0.34$) is slightly lower than the value of 399.0 reported by Borschevsky et al.,10 which was calculated using the CCSD(T) method with the Fægri basis set.65 This minor discrepancy could be due to differences in the basis set and fitting procedures used in this study. Further details can be found in the Supporting Information.
Figure 1(a) demonstrates that non-relativistic polarizabilities increase monotonically with atomic numbers, starting from Na. In contrast, the relativistic data increase with increasing atomic numbers from Na to Cs, but decrease from Cs to Fr, with the value for Fr being close to that of Rb. This trend confirms the irregular progression of values with atomic numbers for Group 1 elements, driven by the relativistic contraction of the valence $s$ shell.61, 66 The polarizability of Na is slightly lower than that of Li due to reduced screening of the nuclear charge by the $p$ electrons in Na, resulting in more compact valence shells and lower polarizabilities.
Figure 1(b) indicates that the scalar-relativistic effect is more pronounced than the SOC effect, leading to a roughly $Z^{2}$ increase in relativistic effects on dipole polarizabilities due to the scalar-relativistic contribution. The SOC contribution to dipole polarizabilities for Group 1 atoms is small; generally, the SOC effect slightly increases the polarizabilities of Group 1 elements, except for Fr, where the SOC effect reduces dipole polarizability by nearly 2. This observation aligns with previous findings.61, 66
4.2 Group 2 elements
As shown in LABEL:tab:dipole_summary, the DC values are in good agreement with the recommended values, except for beryllium (Be, $Z=4$) and Sr, where the DC values slightly exceed the recommended upper bounds. However, the DC result (37.78) for Be using CCSD(T) aligns well with recent theoretical values ($37.73\pm 0.04$ 67, 68 and 37.787 69) from the RCICP method, as well as the CI+MBPT2 result ($37.76\pm 0.22$) based on experimental data.70 In addition, the recommended value ($197.2\pm 0.2$) for Sr in Ref. 1 is based on computational results from Refs. 70, 71, where a hybrid relativistic CI and MBPT method is used. This implies that the discrepancy may arise from higher-order correlation effects included in this work.
Figure 2(a) displays the relationship between dipole polarizabilities and atomic numbers for Group 2 elements. It reveals that non-relativistic polarizabilities increase monotonically with atomic numbers. However, the trend for relativistic polarizabilities resembles the non-relativistic case for lighter atoms but is lower for heavier elements, particularly for Sr and barium (Ba, $Z=56$), where deviations exceed 11 and 42, respectively. Moreover, a considerable decrease in relativistic polarizabilities is observed from Ba to radium (Ra, $Z=88$).
Figure 2: Same as Figure 1, but for Group 2 elements.
4.3 Group 13 elements
As shown in LABEL:tab:dipole_summary, both DC-CC and DC-CI values in this work align well with the recommended theoretical values presented in Ref. 1, except for boron (B, $Z=5$) and Nh. It should be noted that the DC-CC data for B is not available in the present work due to convergence issues encountered with DHF using a DC Hamiltonian. Therefore, the most accurate results are DC-CC results for all Group 13 elements except for B, for which the SR value is used as the most accurate result. For Al, the best DC-CC result is obtained by including 11 outer-shell electrons and 223 virtual shells for electron correlation instead of all electrons and virtual shells, as shown in Table S3. Furthermore, all component-resolved values for Group 13 atoms are consistent with the results computed in Ref. 14. While an earlier study has shown that virtual cutoff values of 10 are sufficient for DC calculations,14 I opted for a tight cutoff of 25 in this work. This choice stems from the confidence in earlier findings and the anticipation that a higher cutoff might ensure augmented accuracy.
For Nh, the computed DC-CC dipole polarizability of $32.89\pm 0.18$ exceeds the recommended value of $29.2\pm 2$. Prior studies using FS-CCSD (29.85) and SD+CI (28.8) reported lower values. Differences arise from the larger basis set and fewer correlated electrons in this work. Validation tests, including two-component DC-CC calculations and a larger basis set, confirm the convergence of the basis set and number of correlated electrons. The computed DC-CI value of $28.21$ agrees with prior SD+CI results. This suggests an updated recommended value for Nh based on the CCSD(T) result is warranted. Further details can be found in the Supporting Information.
Figure 3(a) displays the relationship between dipole polarizabilities and atomic numbers for Group 13 elements. Also, the $M_{L}$ component polarizabilities at the spin-free-relativistic level are presented. In general, polarizabilities increase with atomic numbers at the non-relativistic level, with the exception of Ga, which has a smaller screening of the nuclear charge by $d$ electrons, resulting in more compact valence shells and a reduced polarizability as reported in the previous study.14 From the component-resolved results in Figure 3(a), the decrease in polarizability is mainly caused by the $M_{L}=\pm 1$ state. This effect is also observed in the comparison between In and Tl, but it is smaller than in the Ga case at the non-relativistic level. The difference in $M_{L}=0$ states of In and Tl eliminates the increase in the corresponding $M_{L}=\pm 1$ components, resulting in a polarizability value for Tl that is close to that of In in non-relativistic calculations. However, at the scalar-relativistic level, this change cannot be attributed solely to $f$ electrons screening due to relativity. In the Dirac picture, the SOC effect significantly reduces the polarizabilities of Tl and Nh. As a result, the polarizability of Tl is close to that of Ga, and the result for Nh is lower than Al. This finding is consistent with previous work in Ref. 14.
Both relativistic effects and electron correlation have a reducing effect on the polarizabilities of Group 13 elements, as seen in Figures 3(b) and (c). The DC relativistic effects ($\Delta\alpha_{r}^{\text{DC}}$) roughly increase with atomic numbers to the power of 2, as well as through the SOC effect, which is the dominant factor compared to scalar-relativistic effects ($\Delta\alpha_{r}^{\text{SR}}$). An exception is Al, whose DC-CC calculation is not accurate. The smaller difference between $\bar{\alpha}_{J=3/2}$ and $\alpha_{J=1/2}$ for B and Al in LABEL:tab:dipole_group_13 also suggests that SOC contribution can be disregarded entirely, which is consistent with previous research presented in Ref. 14.
4.4 Group 14 elements
As shown in LABEL:tab:dipole_summary, all DC results are in good agreement with the recommended values in Ref. 1 except for carbon (C, $Z=6$). Additionally, the DC value (11.73) of C is slightly larger than the upper bound of the result ($11.26\pm 0.20$) reported in Ref. 72. The difference could be attributed to the different basis sets and the number of data points computed. The basis set ($19s11p6d4f2g$) is used in this work compared to the ($13s7p4d3f2g$) used in Ref. 72. In addition, energy with an external electric field of $+0.0005$ a.u. is also added in this work compared to Ref. 72. Moreover, the difference between SR and DC values, as shown in LABEL:tab:dipole_summary, could be attributable to the small SOC effect for carbon. This suggests that $LS$ coupling may provide a better description than $jj$ coupling. The calculated SR polarizability of C (11.61) is close to the DK value (11.70) in Ref. 72 and the recent scalar-relativistic CCSD(T) result (11.524).73 Due to the small relativistic effects, numerous NR values ($11.67\pm 0.07$ 37, 11.63 74, $11.64\pm 0.15$ 75, 11.683 76, 11.324 76) computed by the CCSD(T) method have been reported in the literature. The computed NR value (11.66) in this work is in good agreement with these values, with the largest error being less than 2%.
The relationship between dipole polarizabilities and atomic numbers for Group 14 elements is depicted in Figure 4(a). In general, all dipole polarizabilities increase with atomic numbers except for atoms from tin (Sn, $Z=50$) to flerovium (Fl, $Z=114$) due to the SOC effect. Figure 4(b) shows that $\Delta\alpha_{r}^{\text{SOC}}$ and $\Delta\alpha_{r}^{\text{DC}}$ increase approximately as $Z^{2}$ with atomic numbers, while $\Delta\alpha_{r}^{\text{SR}}$ increases linearly as $Z$ with respect to atomic numbers. Moreover, the main difference between $\alpha^{\text{NR}}$ and $\alpha^{\text{SR}}$ is from the $M_{L}=0$ component, and the difference increases with atomic numbers, especially lead (Pb, $Z=82$) and Fl. For lighter atoms C, silicon (Si, $Z=14$), and germanium (Ge, $Z=32$), all results are very close, suggesting that relativistic effects are negligible.
As shown in Figure 4(c), electron correlation reduces the dipole polarizabilities except for the DC value of Fl, which agrees with the previous findings in Ref. 72. The large deviation between $\Delta\alpha_{c}^{\text{SR}}$ and $\Delta\alpha_{c}^{\text{DC}}$ could be attributed to the fact that the full-relativistic DC DHF wave function is not a proper reference wave function, leading to considerable correlation energy for Si, Ge, and Sn.
Figure 4: Same as Figure 3, but for Group 14 elements.
4.5 Group 15 elements
As shown in LABEL:tab:dipole_summary, all DC, SR, and NR polarizabilities of the ground state for Group 15 elements are in excellent agreement with the recommended values reported in Ref. 1, except for the DC value of As (28.98), which is slightly smaller than the lower bound of the corresponding recommended value ($30\pm 1$). This implies that the relativistic effects are small for all Group 15 elements except for Mc, where a large uncertainty is used. Therefore, in this work, the most accurate calculations for nitrogen (N, $Z=7$), phosphorus (P, $Z=15$), and arsenic (As, $Z=33$) are SR CCSD(T), while DC MRCISD is used for antimony (Sb, $Z=51$), bismuth (Bi, $Z=83$), and Mc due to the increasing SOC contribution to polarizabilities.
There are no DC values available for N, P, and As yet, based on Ref. 1, 19. For Sb, only one DC value ($45\pm 11$) is reported by Doolen with relativistic local density approximation (LDA) in linear response theory.77 Additionally, this method is also used to compute the DC value ($50\pm 12$) of Bi. The most accurate DC values for Bi and Mc in the literature are reported by Dzuba et al. using SD+CI methods. It should be noted that the DC values for Bi and Mc in this work are 46.57 and 68.64, respectively, which are in excellent agreement with the SD+CI values (44.62 and 70.5).16
Figure 5: Same as Figure 1, but for Group 15 elements.
4.6 Group 16 elements
As shown in LABEL:tab:dipole_summary, the DC-CI results are in agreement with the recommended values presented in Ref. 1, except for sulfur (S, $Z=16$) and selenium (Se, $Z=34$). Considering the small relativistic effects for light elements, the most accurate calculations for oxygen (O, $Z=8$), S, and Se in this work are SR CCSD(T). In contrast, DC MRCISD is used for tellurium (Te, $Z=52$), polonium (Po, $Z=84$), and Lv due to the increasing SOC contribution to polarizabilities.
Figure 6: Same as Figure 3, but for Group 16 elements.
4.7 Group 17 elements
As shown in LABEL:tab:dipole_summary, the DC values align well with the recommended values, except for the light atoms, F and chlorine (Cl, $Z=17$). Given the small relativistic effects in light elements, the most accurate calculations for F, Cl, and Br in this work are SR CCSD(T). In contrast, DC MRCISD is used for I, At, and Ts due to the increasing SOC contribution to polarizabilities. As expected, the polarizabilities of states $J=\frac{3}{2},M_{J}=\frac{1}{2}$ are found to be smaller than those of the state $J=\frac{3}{2},M_{J}=\frac{3}{2}$. Moreover, the values for $J=\frac{3}{2},M_{J}=\frac{3}{2}$ are also slightly larger than those for $J=\frac{1}{2},M_{J}=\frac{1}{2}$ across all elements except for Ts, which displays an anomalously high polarizability, nearly four times larger than that of the $\alpha_{3/2,1/2}$ state. This deviation can be attributed to the heavily relativistic contraction of the $\phi\left(\frac{1}{2},\frac{1}{2}\right)$ spinors, which are less polarizable and more occupied in the $J=\frac{3}{2}$ states as reported in Ref. 17. The small values of $\alpha_{J=1/2}-\bar{\alpha}_{J=3/2}$ and $\bar{\alpha}_{M_{L}}-\bar{\alpha}_{3/2}$ suggest that the SOC effect can be neglected in the calculation of dipole polarizabilities for Group 17 atoms, except for Ts, where the SOC effects are particularly significant.
Figure 7: Same as Figure 3, but for Group 17 elements.
4.8 Group 18 elements
@@
As shown in LABEL:tab:dipole_summary, the DC results for the heavy atoms, including krypton (Kr, $Z=36$), xenon (Xe, $Z=54$), radon (Rn, $Z=86$), and Og, show excellent agreement with the reference values in Ref. 1. For the lighter atoms, like Ne, and argon (Ar, $Z=18$), the polarizabilities calculated in this study are slightly lower than the recommended values as shown in LABEL:tab:dipole_summary. Moreover, the NR, SR, and DC values are 2.57 for Ne in this work, which is lower than experimental 78, 79, 9, 80 and theoretical values,81, 59, 82, 75, 83, 84 i.e., 2.66. Additionally, the SR and DC values are 11.02 for Ar in this work, which is lower than experimental 85, 86, 79, 59, 80 and theoretical values,59, 87, 88, 89 i.e., 11.08.
@@
Figure 8: Same as Figure 1, but for Group 18 elements.
@@ 5 Summary
    The atomic dipole polarizabilities obtained by fitting Eq. (15) are also included. A summary table of the most accurate results for all main groups is provided.
@@
    The reference atomic dipole polarizabilities from the literature, reported in Refs. 1, 19, are sorted by publication year for each group. A detailed comparison between the computational values in this work and the data from the literature is provided for each atom.
  4. 4.
    A detailed uncertainty estimation is provided for each atom.